Hartley Company sold $900,000, 10-year, 9% bonds on January 1 for $885,000. Interest is to be paid annually on January 1. If the issuing corporation uses the straight-line method to amortize discounts and premiums on bonds payable, the amount of bond interest expense to be recognized in the year issued is

Group of answer choices

$90,000.

$81,000.

$82,500.

$88,500.

$79,500.
Answer #1

Interest payable = 900,000 * 9% = 81,000

Bond discount amortization

= (900,000-885,000)/10 = 1500

Interest expense to be recognized

= 81,000 + 1500

= 82,500